# Safe and Generalized end-to-end Autonomous Driving System with Reinforcement Learning and Demonstrations

### v3

@@ 4.1.1 Traning environment
The training and evaluation of our entire autonomous driving perception-decision-making joint system are conducted on the CARLA(Dosovitskiy et al., 2017). In order to comprehensively evaluate the performance of our SGADS, we utilized six maps in CARLA ranging from Town01 to Town06, as illustrated in Figure 5. Town01, a small, simple town with a river and several bridges. Town02, a small simple town with a mixture of residential and commercial buildings. Town04, a small town embedded in the mountains with a special infinite highway. Town05, squared-grid town with cross junctions and a bridge. Town06, long many lane highways with many highway entrances and exits. Particularly, Town03 is the most complex town with a 5-lane junction, a roundabout, unevenness, a tunnel, and more. To evaluate more realistically, we adopted the approach of randomly generating driving routes, resetting the coordinates of all vehicles in the map, and randomizing the driving scenarios in each episode.
@@
4.1.2 Measure Safety
In this experimental environment setting, high safety is defined as the conditions in which an intelligent vehicle in its environment can follow a randomly predetermined route without colliding with surrounding vehicles, avoid major deviations from a randomly predetermined trajectory, avoid stopping on the road, avoid colliding with a roadside wall or guardrail, and successfully navigate a complex intersection or roundabout without any collision. If the intelligent vehicle fails to comply with any of the above regulatory constraints, it will be penalized, resulting in the termination of the current episode in our setting. In addition, when an episode ends, all scenes and vehicles are forcibly and randomly reset, resulting in the termination of the episode. Therefore, we use the Average Driving Distance (Avg Dis) to measure safety, indicating the average distance that an intelligent vehicle can safely drive in multiple episodes. When the intelligent vehicle drives a longer driving distance within the episode, it indicates that its safety performance is higher. Simultaneously, as the episode length increases and our testing extends, the intelligent vehicle is more likely to have a longer safe driving distance. Additionally, we evaluate the ability to safely travel the maximum distance over multiple episodes using Max Driving Distance(Max Dis), calculated as the maximum value of the driving distance across multiple episodes.
4.2 Human expert collects teaching datasets
4.3 Evaluate the prediction of driving trajectories
In order to accurately evaluate our model prediction of driving behavior for intelligent vehicle in the imagination space, this problem can be viewed as a special POMDP problem with the reward value maintained at 0. Then, two scenarios of original sensor inputs for intelligent vehicle at time steps 1-15 are randomly sampled as shown in Figure 4, followed by further allowing our model to predict the driving behaviors of intelligent vehicle at time steps 16-30 under an imagination space length of $H=15$. The result shows that our model can accurately predict future driving trajectories.
Figure 7: All training for SGADS was completed in Town03, followed by separate evaluations from Town01 to Town06. In all evaluations, the calculations for Avg Dis and Max Dis are based on 5 episodes. Due to the smaller size of the map in town02, the number of vehicles is limited to 50, while in other maps, the number of vehicles is set to 200.
4.4 Evaluate Sampling efficiency
In order to evaluate the performance of our autonomous driving system more effectively, we have conducted various comparisons with existing methods such as DDPG(Lillicrap et al., 2015), SAC(Haarnoja et al., 2018), TD3(Fujimoto et al., 2018), DQN(Mnih et al., 2015) and Latent_SAC(Chen et al., 2021). These methods achieve their best performance when using lidar as input: As shown in Table 1, compared to other methods, our SGADS achieves at least a 30 times and 10 times efficiency improvement in training time when reaching baselines of 50 meters and 150 meters, respectively, with Lidar as input. Particularly, when the baseline is set at 200 meters and the training time exceeds 250 hours, only our model achieves this baseline. Additionally, by exploring the input as Lidar_ng, efficiency of the model is further enhanced compared to Lidar. This indicates that our SGADS can effectively reduce training costs and improve sampling efficiency compared to existing methods.
4.5 Evaluate Safety
To comprehensively evaluate safety, we chose the most complex urban map, Town03. Our SGADS is initially trained in Town03 and subsequently evaluated in Town03 under random and complex scenarios. In random scenarios, the ego vehicle and around vehicles can appear at any location on the map. In fixed scenarios, they are restricted to random appearances within a specific scene range. The difference between the two lies in the range of random selection, but both follow the randomization mechanisms of CARLA in every episode during training and evaluation.
As shown in Figure 7, the first row from left to right illustrates the three scenarios we evaluated in Town03, including random driving scenarios, fixed complex roundabouts, and fixed intersections. The experimental results show that SGADS exhibits outstanding Avg Dis and Max Dis in both random driving scenarios and complex scenarios, significantly surpassing existing methods. This further emphasizes that our approach has the potential to significantly enhance the safety of intelligent vehicles.
Additionally, to clarify the sources of safety benefits, we decomposed SGADS into three components, variational inference based on normalizing flows (Base), safety constraints (SC), and behavior cloning (BC). We then conducted evaluations using three comparison settings, Base, Base+SC, and Base+SC+BC. This is illustrated in the first row of evaluations in Figure 7 for the ablation evaluation in Town03. Under the input conditions of lidar or Lidar_noground, the safety benefits of SGADS primarily result from the design of the SC component, while the BC component can further provide additional benefits.
4.6 Evaluate Generalization
To evaluate the generalization performance of SGADS, in each episode of the experiment, scenes are randomly set, driving routes are generated randomly, and environmental vehicles are randomly generated, ensuring the complexity and diversity of the evaluation. As shown in the second row of Figure 7, we initially train SGADS and other methods in Town03 and evaluate respectively in Town01, Town02, Town04, Town05, and Town06. The experimental results demonstrate that SGADS exhibits significantly superior Avg Dis and Max Dis compared to other methods, and the required training time is relatively short, confirming its efficient training performance.
Additionally, despite the relatively small size of the Town02 map, resulting in a smaller upper limit for driving distance, SGADS still outperforms other methods in terms of Avg Dis, Max Dis, and training efficiency. This indicates that SGADS exhibits outstanding generalization capabilities, primarily benefiting from the integration of various aspects, including realistic random mechanisms close to the real world in CARLA, trajectory prediction based on normalizing flows, safety constraints, and the fusion of human expert demonstrations.
@@ 4.6 Evaluate Generalization
Table 2: Evaluate the performance of intelligent vehicle under multiple input images in town03, training steps=100k. Avg Dis is based on 5 episodes, the number of vehicles is 200.
Multiple input Avg dis(m)
@@
4.7 Multiple types of input images
We compare various input image types for evaluating the performance of intelligent vehicle based on BC of the SGADS, as shown in Table 2. The representation of multiple inputs is described in the section Appendix A.1 Avg Dis of lidar_noground reaches the highest value compared with all other input types. This is because lidar_noground removes a large amount of redundant information, reduces the difficulty of world model understanding environment semantics, and also involves stationary status of intelligent vehicle in experiment. The results show that the lidar_noground input is relatively optimal. However, it is worth noting that the effects of these 12 different input types are relatively small, with the Avg Dis e only varying between 20 and 40 meters. This shows that different data types have a minimal impact on the safety performance of intelligent vehicles.
5 Conclusion
In order to address the challenges of low sample efficiency, low safety, limited generalization, and accurate prediction of future driving trajectories in autonomous driving systems, we propose a safe, efficient, and generalizable end-to-end autonomous driving system (SGADS). Experimental results demonstrate that compared to existing methods, the safety constraints in SGADS significantly enhance the safety of intelligent vehicles, and demonstrations further improve safety and sample efficiency. Variational inference based on normalizing flow accurately predicts future driving trajectories, providing a precondition for optimal decision making. Finally, the fusion of these three modules greatly enhances the generalization capabilities of intelligent vehicles, significantly surpassing existing methods. Additionally, we observe that despite differences in input data types, these variances have minimal impact on the safety. We contribute a demonstration dataset by having human experts drive using the G29 steering wheel. However, there is still space for improvement in aligning our SGADS system with real-world safety and generalization requirements.
Appendix A Appendix
A.1 Multiple types of input images
The 12 types of input data we designed are mainly categorized into single-modal and single-image input, single-image and multimodal fusion, and multiple images and multimodal fusion.
@@ A.1 Multiple types of input images
A.2 Reward function
We use the following reward function in our experiments:
$\text{R}=200\cdot r_{c}+200\cdot r_{ft}+50\cdot r_{lt}+2\cdot r_{sc}+v_{lon}+10\cdot r_{f}+r_{o}-5\alpha^{2}+0.2\cdot r_{lat}-0.1$ (12)
where $r_{c}$ is the reward related to collision, which is set to -1 if the ego vehicle collides and 00 otherwise. $v_{lon}$ is the longitudinal speed of the ego vehicle. $r_{f}$ is the reward related to running too fast, which is set to $-1$ if it exceeds the desired speed (8 m/s here) and 00 otherwise. $r_{o}$ is set to $-1$ if the ego vehicle runs out of the lane, and 00 otherwise. $\alpha$ is the steering angle of the ego vehicle in radians. $r_{lat}$ is the reward related to lateral acceleration, which is calculated by $r_{lat}=-|\alpha|\cdot v_{lon}^{2}$. The last constant term is added to prevent the ego vehicle from standing still. $r_{ft}$ represents the time to collision in the forward direction. $r_{lt}$ represents the time to collision in the lateral direction. $r_{sc}$ represents the smoothness constraint.
A.3 Hyperparameter settings
$\mathcal{M}_{model}$, the KL regularizer is clipped below 3.0 free nats for imagination range $H=15$ using the same trajectories for updating action and value models separately with $\lambda=0.99$ and $\lambda=0.95$, while $k=1.5$. The size of all our trainig and evaluating images is $128\times 128\times 3$. For Algorithm 1, a random seed $S=5$ is used to collect datasets for the $ego$ vehicle before updating the model every $C=100$ steps during training process.
@@ A.3 Hyperparameter settings
SGADS 32 32 10 1
@@
SGADS $1\times 10^{-3}$ $8\times 10^{-5}$ $8\times 10^{-5}$
A.4 The world model reconstructs the input images from the original sensors
@@
Figure 8: Randomly sampled frames to reconstruct the input images from the original sensors of SGADS on 8 types of input. For each type of image, first row: original sensor inputs. Second row: reconstructed images.
A.5 Algorithm
we propose the framework of safe and generalized end-to-end autonomous driving system with reinforcement learning and demonstrations is presented in Algorithm 1.
Algorithm 1 SGADS
Input: batch size $B$, batch length $L$, imagination horizon $H$, step $T$, trajectory $\tau$, number of trajectories $N$, initialize datasets $\mathcal{D}$ with seeds.
1:  The first stage of trainingCollect $\mathcal{D}_{expert}$ via the G29 steering.Randomly initialize parameters $\theta,\phi,\psi,\eta,\omega$.
2:  while $\mathcal{M}_{expert}$ does not converge do
3:     for $c=1,2,\ldots,C$ do
4:        Randomly sample $B$ sequences $\{(a_{t}^{\prime},o_{t}^{\prime},r_{t}^{\prime})\}_{t=k}^{k+L}$ from $\mathcal{D}_{expert}$.
5:        Calculating the states $p_{\eta}(s_{t}^{\prime}|s_{t-1}^{\prime},a_{t-1}^{\prime},o_{t}^{\prime})$.
6:        Update parameters $\phi,\eta$.
7:     end for
8:  end while
9:  The second stage of trainingInitialize all neural network parameters using $\mathcal{M}_{expert}$.
10:  while $\mathcal{M}_{model}$ does not converge do
11:     for $c=1,2,\ldots,C$ do
12:        Ramdomly sample $B$ sequences from $\mathcal{D}$.
13:        Infer belief state $s_{t}\sim q_{K}(s_{t}|s_{t-1},a_{t-1},o_{t})$.
14:        for $i=1,2,\ldots,N$ do
15:           Rollout imaginary trajectories ${(s_{i,\tau},a_{i,\tau})}_{\tau=t}^{t+H}$ with belief transition model.
16:        end for
17:        Update $\theta,\phi,\psi,\xi,\eta,\omega$.
18:     end for
19:     for $t=1,2,\ldots,T$ do
20:        Compute $s_{t}\sim q_{K}(s_{t}|s_{t-1},a_{t-1},o_{t})$ from history.
21:        Compute $a_{t}\sim\pi(a_{t}|s_{t})$ with action model.
22:        $r_{t},o_{t+1}\leftarrow$ TTC and CSS.
23:     end for
24:     Update $\mathcal{D}\leftarrow\mathcal{D}\cup\{(a_{t},o_{t+1},r_{t})_{t=1}^{T}\}$
25:  end while
A.6 Further comparisons were made regarding safety and generalization
In the existing baseline methods, the default number of environment vehicles is set to 100. We also conducted comparisons in terms of generalization and safety as shown in Figure 9 and 10. The experiments indicate that our model significantly outperforms existing methods.
Figure 9: The model is trained in Town03 and then evaluated for generalization in Town01, Town04 and Town05.
Figure 10: The model is trained in Town03 and then evaluated for safety in roundabout of the Town03, evluated for generalization in Town06.
A.7 More results regarding predictions of future driving trajectories
@@
Figure 11: We randomly sampled input images, and then SGADS was used to make predictions
@@
Figure 12: We randomly sampled input images, and then SGADS was used to make predictions
@@
Figure 13: We randomly sampled input images, and then SGADS was used to make predictions
@@
Figure 14: We randomly sampled input images, and then SGADS was used to make predictions
Figure 15: We randomly sampled input images, and then SGADS was used to make predictions.
Figure 16: We randomly sampled input images, and then SGADS was used to make predictions
Figure 17: We randomly sampled input images, and then SGADS was used to make predictions
Figure 18: We randomly sampled input images, and then SGADS was used to make predictions

### v8

@@ A.1 Training CARLA maps
In order to comprehensively evaluate the performance of our EGADS, we utilized four maps in CARLA, Town03, Town04, Town05 and Town06 as shown in Figure 4. Town04, a small town embedded in the mountains with a special infinite highway. Town05, squared-grid town with cross junctions and a bridge. Town06, long many lane highways with many highway entrances and exits. Particularly, Town03 is the most complex town with a 5-lane junction, a roundabout, unevenness, a tunnel, and more.
A.2 Collect expert datasets
A.3 Multiple types of input images
Figure 4: The road networks of the CARLA include routes for Town03, Town04, Town05, and Town06
@@ A.3 Multiple types of input images
The 12 types of input data we designed are mainly categorized into single-modal and single-image input, single-image and multimodal fusion, and multiple images and multimodal fusion, as shown in Figure 6. We compare various input image types for evaluating the performance of NFRL, as shown in Table 7. ASD of lidar_noground reaches the highest value compared with all other input types. This is because lidar_noground removes a large amount of redundant information, reduces the difficulty of world model understanding environment semantics, and also involves stationary status of intelligent vehicle in experiment. The results show that the lidar_noground input is relatively optimal. However, it is worth noting that the effects of these 12 different input types are relatively small, with the ASD only varying between 20 and 40 meters. This shows that different data types have a minimal impact on the safety performance of intelligent vehicles.
@@
Table 7: Evaluate the performance of NFRL under multiple input images in town03, training steps=100k. ASD is based on 5 episodes, the number of vehicles is 200.
Multiple input ASD (m)
@@
Figure 6: Multiple types of input images
A.4 Reward function
We use the following reward function $f_{2}$ in our experiments: where $r_{c}$ is the reward related to collision, which is set to -1 if the ego vehicle collides and 0 otherwise. $v_{lon}$ is the longitudinal speed of the ego vehicle. $r_{f}$ is the reward related to running too fast, which is set to 1-1 if it exceeds the desired speed (8 m/s here) and 0 otherwise. $r_{o}$ is set to 1-1 if the ego vehicle runs out of the lane, and 0 otherwise. $\alpha$ is the steering angle of the ego vehicle in radians. $r_{lat}$ is the reward related to lateral acceleration, which is calculated by $r_{lat}=-|\alpha|\cdot v_{lon}^{2}$. The last constant term is added to prevent the ego vehicle from standing still. $r_{ft}$ represents the time to collision in the forward direction, and if it is an autonomous vehicle and the time to collision with surrounding vehicles is below the safety threshold, this term is set to -1. $r_{lt}$ represents the time to collision in the lateral direction, and if it is an autonomous vehicle and the time to collision with surrounding vehicles is below the safety threshold, this term is set to -1. $r_{sc}$ represents the smoothness constraint, and if the actual steering angle of the autonomous vehicle differs significantly from the predicted steering angle by the model, exceeding a set empirical constant, this term is set to -1.
[…]
where the reward function $f_{1}$ is proposed by [6].
A.5 Measure performance metrics
We use multiple key metrics to evaluate the performance of autonomous driving models in various driving scenarios. Outlane Rate (OR): the rate at which the vehicle deviates from its designated lane. This metric evaluates the ability of modes to maintain proper lane discipline. Episode Completion Rate (ER): the percentage of driving tasks or episodes that the vehicle successfully completes. Higher completion rates indicate better task performance. Average Safe Driving Distance (ASD): the average distance driven without incidents, such as collisions or off-road events. This metric highlights the capability to drive safely over extended periods. Average Return (AR): A metric that measures the cumulative reward collected by the vehicle during its driving tasks, often reflecting both task performance and adherence to safety guidelines. Driving Score (DS): A comprehensive metric that reflects the overall performance of the vehicle in terms of safety, efficiency, and compliance with traffic rules.
$OR=\frac{N_{\text{off\_road\_events}}}{N_{\text{total\_episodes}}},ER=\frac{N_{\text{completed\_steps}}}{N_{\text{total\_steps}}},AR=\frac{\sum_{i=1}^{N_{\text{episodes}}}\text{rewards}_{i}}{N_{\text{total\_episodes}}}$ (15)
$ASD=\frac{\sum_{i=1}^{N_{\text{episodes}}}\text{distance}_{i}}{N_{\text{total\_episodes}}},DS=ER\times AR$ (16)
Where $N_{\text{total\_episodes}}$ is the total number of episodes in the test. Where $N_{\text{off\_road\_events}}$ is the number of times the vehicle went off-road, and $N_{\text{total\_steps}}$ is the total number of episodes. Where $\text{distance}_{i}$ is the distance driven during the $i$-th safe driving episode, and $N_{\text{safe\_episodes}}$ is the number of episodes without incidents (such as collisions or off-road events). Where $N_{\text{completed\_steps}}$ is the number of successfully completed steps, and $N_{\text{total\_steps}}$ is the total number of steps in the episode. Where $AR$ is the average reward $f$ collected during the episode.
@@ A.5 Measure performance metrics
Dreamer 256 32 5 2
NFRL 32 32 10 1
NFRL+SC 32 32 10 1
BC+Demo 32 32 10 1
NFRL+SC+Demo 32 32 10 1
@@
Dreamer $1\times 10^{-3}$ $8\times 10^{-5}$ $8\times 10^{-5}$
NFRL $1\times 10^{-3}$ $8\times 10^{-5}$ $8\times 10^{-5}$
NFRL+SC $1\times 10^{-3}$ $8\times 10^{-5}$ $8\times 10^{-5}$
BC+Demo $1\times 10^{-3}$ $8\times 10^{-5}$ $8\times 10^{-5}$
NFRL+SC+Demo $1\times 10^{-3}$ $8\times 10^{-5}$ $8\times 10^{-5}$
A.6 Hyperparameter settings
$\mathcal{M}_{model}$, the KL regularizer is clipped below 3.0 free nats for imagination range $H=15$ using the same trajectories for updating action and value models separately with $\lambda=0.99$ and $\lambda=0.95$, while $k=1.5$. The size of all our trainig and evaluating images is $128\times 128\times 3$. A random seed $S=5$ is used to collect datasets for the egoego vehicle before updating the model every $C=100$ steps during training process. We present the hyperparameter settings for the methods involved in our experiments as shown in Table8 and Table 9.
A.7 The world model reconstructs the input images from the original sensors
@@
Figure 7: Randomly sampled frames to reconstruct the input images from the original sensors of EGADS on 8 types of input. For each type of image, first row: original sensor inputs. Second row: reconstructed images.
A.8 More results regarding predictions of future driving trajectories
@@
Figure 8: We randomly sampled input images, and then EGADS was used to make predictions
@@
Figure 9: We randomly sampled input images, and then EGADS was used to make predictions
@@
Figure 10: We randomly sampled input images, and then EGADS was used to make predictions
@@
Figure 11: We randomly sampled input images, and then EGADS was used to make predictions